# Discrete time crystal for periodic-field sensing with quantum-enhanced precision

### v1

@@ VI Experimental proposal
where $n^{\mu}_{j}{=}c^{\mu\dagger}_{j}c^{\mu}_{j}$ counts the particle numbers at site $j$ in chain $\mu{=}a,b$, and $c^{\mu\dagger}_{j}$ ($c^{\mu}_{j}$) is the fermionic creation (annihilation) operator. In this notation, the AC magnetic field can be rewritten as
@@
Certain fermionic atoms, such as dysprosium (Dy) and erbium (Er), exhibit significant magnetic dipole moments. When cooled to ultracold temperatures and loaded into an optical lattice, these atoms can demonstrate behaviors governed by magnetic dipolar interactions. The magnetic dipole-dipole interactions, which are long-range and anisotropic, take the following form
@@
where $r_{ab}{=}\sqrt{y^{2}+|j_{a}-j_{b}|^{2}x^{2}}$. Setting $y{\gg}x$, leads to $V_{j_{a}{=}j_{b}}{\rightarrow}0$ and $V_{j_{a}{\neq}j_{b}}{\rightarrow}0$ and allows tunneling for duration $t_{2}{\simeq}\hbar\pi y^{3}(1{-}\varepsilon){/}2C_{dd}$.
@@
This result shows that, although, for infinitesimal $h$ and $y{\gg}x$, one can recover Eq. (19), larger values of $h$ affect the DTC nature through inducing site-dependent dipole-dipole interactions.
VII Conclusion
In this paper, we propose a DTC in a disorder-free system for sensing a gradient periodic-field with quantum-enhanced sensitivity. Our analysis results in $\mathcal{F}_{Q}{\leqslant}4n^{2}L^{2}(L{-}1)^{2}$ as the ultimate precision limit in the estimation of the gradient periodic-field. Our results show that a many-body probe initialized in a product state can achieve ultimate precision, namely $\mathcal{F}_{Q}{\propto}n^{2}L^{4}$, provided that it operates in the DTC phase. This enhancement is due to the resonance dynamics between two GHZ-type eigenstates of the system. Remarkably, this quantum enhancement shows robustness against different imperfections in the sensing protocol. In addition, we discuss the implementation of our probe in an optical lattice setup. Our results open avenues for harnessing inherent entanglement in interacting, non-equilibrium quantum many-body systems for enhanced metrology.
VIII Acknowledgment
AB acknowledges support from National Natural Science Foundation of China (Grants Nos. $12050410253$, $92065115$, and $12274059$) and the Ministry of Science and Technology of China (Grant No. QNJ2021167001202116700120211670012021167001L.

### v2

@@ VIII Experimental proposals
where $n^{\mu}_{j}{=}c^{\mu\dagger}_{j}c^{\mu}_{j}$ counts the particle numbers at site $j$ in chain $\mu{=}a,b$, and $c^{\mu\dagger}_{j}$ ($c^{\mu}_{j}$) is the fermionic creation (annihilation) operator. In this notation, the periodic magnetic field can be rewritten as
@@
Certain fermionic atoms, such as dysprosium (163Dy) and erbium (167Er) exhibit significant magnetic dipole moments as ${\approx}10\mu_{B}$ and ${\approx}7\mu_{B}$, respectively. Here, $\mu_{B}$ is the Bohr magneton. When cooled to ultracold temperatures and loaded into an optical lattice, these atoms can demonstrate behaviors governed by magnetic dipolar interactions. The magnetic dipole-dipole interactions, which are long-range and anisotropic, take the following form
@@
where $r_{ab}{=}\sqrt{y^{2}+|j_{a}-j_{b}|^{2}x^{2}}$. Setting $y{\gg}x$, leads to $V_{j_{a}{=}j_{b}}{\rightarrow}0$ and $V_{j_{a}{\neq}j_{b}}{\rightarrow}0$ and allows tunneling for duration $t_{2}{\simeq}\hbar x^{3}{/}C_{dd}$.
@@
This result shows that, for infinitesimal $h$ and $y{\gg}x$, one can recover Eq. (23) that allows tunneling due to the suppression of the dipole-dipole interaction. However, larger values of $h$ through inducing site-dependent dipole-dipole interactions affect the tunneling between chains and, hence, demolish the DTC nature. For the lattice spacing reported in a typical experiment $x{\simeq}266~\mathrm{nm}$ (chain spacing $y{\simeq}10x$), the nearest-neighbour dipolar coupling produces an interaction energy $V_{j_{a}{=}j_{b}}/(2\pi\hbar)\simeq 60\mathrm{Hz}$ for Dy and $V_{j_{a}{=}j_{b}}/(2\pi\hbar)\simeq 29.4\mathrm{Hz}$ for Er. Using $t_{2}\simeq\hbar x^{3}/C_{dd}$ one obtains $t_{2}\simeq 1/(2\pi\cdot 60)\approx 2.65~\mathrm{ms}$ for Dy and $\simeq 1/(2\pi\cdot 29.4)\approx 5.41~\mathrm{ms}$ for Er. Adopting the binary-quench timing used here ($t_{1}=t_{2}=1/(2J_{z})$ with the energy scale chosen so that $J_{z}/(2\pi)\sim f_{j_{a}{=}j_{b}}$), the drive period is $T\sim 2.65~\mathrm{ms}$ (Dy) and $\sim 5.41~\mathrm{ms}$ (Er), so a single-cycle repetition rate is $1/T\sim 377~\mathrm{Hz}$ (Dy) and $\sim 185~\mathrm{Hz}$ (Er). For typical coherence time $T_{2}=0.1\ \mathrm{s}$ in the dipolar systems one has $\left\lfloor\frac{100}{2.65}\right\rfloor=37$ for Dy and $\left\lfloor\frac{100}{5.41}\right\rfloor=18$ for Er of cycles possible. Therefore, using a DTC sensor that includes $L$ atoms can result in $0.0270/L^{2}~pT/m/\sqrt{\mathrm{Hz}}$ and $0.0555/L^{2}~pT/m/\sqrt{\mathrm{Hz}}$ using Dy and Er, respectively.
IX Conclusion
In this paper, we propose a DTC in a disorder-free system for sensing a gradient periodic-field with quantum-enhanced sensitivity. Our results show that a many-body probe initialized in a product state can achieve ultimate precision, namely $\mathcal{F}_{Q}{\propto}n^{2}L^{4}$, provided that it operates in the DTC phase. The $L^{4}$ scaling of the QFI is due to the resonance dynamics between two GHZ-type eigenstates in the DTC regime, which amplify the response of the system. The coherent accumulation over $n$ Floquet cycles leads to the observed $n^{2}$ time scaling, and the resulting sensing bandwidth is centered at the drive resonance and limited by the coherence time $T_{2}$ of the system. This enhancement is experimentally achievable through measuring the standard projective measurement. Remarkably, this quantum enhancement shows robustness against different imperfections and dephasing noise in the sensing protocol. Our results show that the DTC sensor that operates in the DTC phase for $n{\leq}50$ periodic quench keeps the behavior of the QFI for a periodic-filed with frequency offset $\delta f$ of the DTC’s frequency when $\delta fn{\leq}1{/}2$, for a crosstalk with strength $\eta{<}1$, for an initialization with perturbation parameter $\vartheta{\leq}10^{-2}\pi$, for an imperfect quench of strength $\varepsilon{<}0.2$, and finally for a local dephasing noise with rate $\Gamma{\leq}10^{-3}$. In addition, we discuss the implementation of our probe in an optical lattice setup. Our results open avenues for harnessing inherent entanglement in interacting, non-equilibrium quantum many-body systems for enhanced metrology.
X Acknowledgment
AB acknowledges support from National Natural Science Foundation of China (Grants Nos. $12050410253$, 9206511592065115, and $12274059$).